How are you tutors doing? Thanks again for this wonderful site.
Please explain how to solve this exponents question and similar questions:
If 2^x - 2^(x-2) = 3(2^13), what is the value of x?
Please note that 2^x means 2 raised to the power x and 2^(x-2) means 2 raised to the power (x-2)

Thanks for your time
Tolly

Answer by mukhopadhyay(490)

I think you have mistakenly posted the same problem twice.
2^x - 2^(x-2) = 3(2^13)
=>2^x(1-1/4) = 3(2^13)
=>2^x(3/4) = 3(2^13)
=>2^x = 2^15
=>x=15;
Answer: x belongs to {15}
You have a unique name (in good sense), never heard before.